solve the given system of equations
x+y=5
x-y=1

Answer by stanbon(75887)

x+y=5
x-y=1
------
Add and solve for "x":
2x = 6
x = 3
--------
Solve for "y":
x + y = 5
3 + y = 5
y = 2
----------------
Cheers,
Stan H.
============